directions
find the rate of change for the sitution.
you run 7 miles in one hour and 21 miles in three hours?

Answer by stanbon(75887)

find the rate of change for the sitution.
you run 7 miles in one hour and 21 miles in three hours?
---
You have 2 points (1,7) and (3,21)
------------------------
rate of change = (21-7)/(3-1) = 15/2 = 7 mph
=================================================
Cheers,
Stan H.